OpenStudy (anonymous):

whats the inverse of 3x-3

OpenStudy (saifoo.khan):

y = 3x - 3 Solve for x.

OpenStudy (saifoo.khan):

\[\frac{y + 3 }{3} = x\]Therefore inverse; \[f^{-1}(x) = \frac{x+3}{3}\]